# A Runtime-Adaptive Transformer Neural Network Accelerator on FPGAs

### v1

@@ 5.4. Model for BRAM utilization
Here, BRAM_d is the depth of BRAMs which indicates the number of storage locations (or entries) within a BRAM block. Each location holds a fixed number of bits, defined by the width of BRAM (BRAM_w), and both parameters can vary depending on the platform. Bit_w is the bit precision of the data being stored. BRAM_w = 36 and BRAM_d = 1024 for most FPGAs.
@@ 6. Evaluation and Results
Fig. 10 compares the power consumption (in watts) and power efficiency (throughput per watt, GOPS/W) for various models across different CPUs, GPUs, and our FPGA accelerator. Data for different models and platforms were obtained from cited literature, and we used them to compare the performance of ADAPTOR’s performance on the U55C platform for the same models. Since ADAPTOR is synthesized only once, and power is measured using Vivado’s power estimation tool post-synthesis, the total dynamic power consumption remains constant for all models. The JETSON TX2 GPU (li_ftrans_2020, ) achieves the highest power efficiency for the BERT model, mainly due to the sparse architecture of the algorithm, and also has the lowest overall power consumption. The RTX K5000 GPU (han_hpta_nodate, ) is 1.5$\times$ more power efficient than ADAPTOR for the BERT model, due to compression techniques, but consumes 10$\times$ more power. The i7-8700K CPU is the least power-efficient for BERT (han_hpta_nodate, ). ADAPTOR is 1.2$\times$ and 2.87$\times$ more power efficient than the NVIDIA K80 GPU and i7-8700K CPU, respectively, when running BERT, according to FQ-BERT (liu_hardware_2021, ). A custom encoder with four encoding layers was run on an i5-4460 CPU and an RTX 3060 GPU (yang_efa-trans_2022, ), both of which were 5.1$\times$ and 1.63$\times$ less power efficient than ADAPTOR while also being more power-hungry. Fang et al.(fang_algorithm, ) executed a shallow transformer on an i9-9900X CPU, JETSON NANO GPU, RTX 2080, and RTX 3090 GPUs. Although the JETSON NANO GPU consumed 1.56$\times$ less power than ADAPTOR, the other devices used 14–30$\times$ more power. However, ADAPTOR is 3.7$\times$, 1.28$\times$, 4.4$\times$, and 1.67$\times$ more power efficient than all of them.
@@
Fig. 12 illustrates the roofline model of ADAPTOR showing the peak performance and peak memory bandwidth any application can achieve with it. The Memory Bound (blue dashed line) indicates the highest performance attainable based on the available memory bandwidth, which is 200 kB/s. Any data point positioned to the left of this line signifies that performance is constrained by memory bandwidth. The Compute Bound (red line) represents the maximum performance achievable by the FPGA’s computational resources, limited to 0.053 TOPS. Points below the compute bound line indicate that the computational resources are underutilized. All data points (green, yellow, and purple) fall within the compute and memory bound regions, meaning none of them fully utilize the available resources of ADAPTOR. The yellow square, representing the BERT model with $TS_{MHA}=64$ and $TS_{FFN}=192$, achieves the highest performance, being closest to the compute bound. In contrast, the purple star, representing the shallow transformer model with $TS_{MHA}=64$ and $TS_{FFN}=128$, has the highest operational intensity but the lowest performance.
@@
[…]
Table 1 compares the performance of our accelerator, ADAPTOR, with other FPGA-based accelerators. Each of these accelerators is optimized for specific TNN models, with some designed specifically for sparse computations. TRAC (plagwitz_trac_2022, ) is the only one that automatically generates accelerator code based on the target FPGA and TNN architecture. Since ADAPTOR was synthesized once with fixed hardware resources and bit width, and implemented on a different platform without sparsity, we evaluated metrics such as throughput (GOPS), power consumption, normalized throughput (GOPS per DSP or GOPS per LUT), and power efficiency (GOPS per watt) for a fair comparison. ADAPTOR achieved 1.9$\times$ and 2.25$\times$ higher GOPS compared to the accelerators by Qi et al. in (qi_accommodating_2021, ) and (qi_accelerating_2021, ), respectively, for a shallow transformer. Its normalized throughput was also higher, indicating more efficient DSP and LUT usage without relying on sparsity, whereas Qi et al. applied block balanced pruning and complex block row storage format. Fang et al. (fang_algorithm, ) reported higher throughput, normalized throughput, and power efficiency than ADAPTOR, using HDL design techniques and 75% sparse transformer. Qi et al.’s four-layer transformer encoder (qi_accelerating_2021, ) was 1.7$\times$ slower and 2$\times$ less resource-efficient than ADAPTOR although they applied hierarchical pruning. FTRANS (li_ftrans_2020, ), FQ-BERT (liu_hardware_2021, ), and TRAC (plagwitz_trac_2022, ) used various weight compression techniques to achieve very high GOPS, throughput, and power efficiency for the BERT model. FTRANS used a block-circulant matrix-based weight representation. FQ-BERT compressed BERT by integer and fixed point quantization. However, since their implementations are not fully open source and difficult to interpret, a direct comparison with ADAPTOR is challenging. It is unclear how FQ-BERT achieved such high throughput with low DSP and LUT utilization. FTRANS used more DSPs and LUTs than our design, but how they managed this in HLS without long compilation time and without tiling remains unclear. TRAC (plagwitz_trac_2022, ) also used fewer DSPs and LUTs but reported 3.2$\times$ higher GOPS and 8.4$\times$ higher GOPS/DSP, which is puzzling. None of these accelerators applied tiling or described any partitioning scheme to accelerate large models such as BERT, which seems impractical. Tzanos et al.’s work (tzanos_hardware_2022, ) is more comparable to ADAPTOR, as they applied tiling and utilized more resources to achieve speeds 1.6 times faster with GOPS / DSP similar to ADAPTOR.

### v2

@@ 5.4. Model for BRAM utilization
Here, BRAM_d is the depth of BRAMs, which indicates the number of storage locations (or entries) within a BRAM block. Each location holds a fixed number of bits, defined by the width of BRAM (BRAM_w), and both parameters can vary depending on the platform. Bit_w is the bit precision of the data being stored. BRAM_w = 36 and BRAM_d = 1024 for most FPGAs. Each term in the equation corresponds to an array declared in the HLS code. For instance, the first term represents the number of BRAMs synthesized for 10 arrays of size SL×$\times$ dmodel. The max function in the second term accounts for cases where an array may not fully utilize the 18 kb width of a synthesized BRAM, but at least one 18 kb BRAM will still be allocated. The factor 0.5 is used because the total BRAM count is calculated based on 36 kb BRAMs.
@@ 6. Evaluation and Results
Fig. 10 compares the power consumption (in watts) and power efficiency (throughput per watt, GOPS/W) for various models across different CPUs, GPUs, and our FPGA accelerator. Data for different models and platforms were obtained from cited literature, and we used them to compare the performance of ADAPTOR on the U55C platform for the same models. Since ADAPTOR is synthesized only once, and power is measured using Vivado’s power estimation tool post-synthesis, the total dynamic power consumption remains constant for all models. The JETSON TX2 GPU (li_ftrans_2020, ) achieves the highest power efficiency for the BERT model, mainly due to the sparse architecture of the algorithm, and also has the lowest overall power consumption. The RTX K5000 GPU (han_hpta_nodate, ) is 1.5$\times$ more power efficient than ADAPTOR for the BERT model, due to compression techniques, but consumes 10$\times$ more power. The i7-8700K CPU is the least power-efficient for BERT (han_hpta_nodate, ). ADAPTOR is 1.2$\times$ and 2.87$\times$ more power efficient than the NVIDIA K80 GPU and i7-8700K CPU, respectively, when running BERT, according to FQ-BERT (liu_hardware_2021, ). A custom encoder with four encoding layers was run on an i5-4460 CPU and an RTX 3060 GPU (yang_efa-trans_2022, ), both of which were 5.1$\times$ and 1.63$\times$ less power efficient than ADAPTOR while also being more power-hungry. Fang et al.(fang_algorithm, ) executed a shallow transformer on an i9-9900X CPU, JETSON NANO GPU, RTX 2080, and RTX 3090 GPUs. Although the JETSON NANO GPU consumed 1.56$\times$ less power than ADAPTOR, the other devices used 14–30$\times$ more power. However, ADAPTOR is 3.7$\times$, 1.28$\times$, 4.4$\times$, and 1.67$\times$ more power efficient than all of them.
@@
Fig. 12 presents the roofline model of our accelerator, highlighting its peak performance and memory bandwidth limits. The Memory Bound (blue dashed line) indicates the maximum achievable performance based on the memory bandwidth, which is 103,000 GB/s. Data points to the left of this line are constrained by memory bandwidth. The Compute Bound (red line) represents the peak performance determined by the FPGA’s computational resources, capped at 53 GOP/s. Points below this line indicate underutilization of computational resources. All data points (green, yellow, and purple) fall within the compute and memory-bound regions, meaning none fully utilize the accelerator’s available resources. The yellow square, representing the BERT model with $TS_{MHA}=64$ and $TS_{FFN}=192$, achieves the highest performance, being closest to the compute bound. In contrast, the purple star, corresponding to the shallow transformer model with $TS_{MHA}=64$ and $TS_{FFN}=128$, exhibits the highest operational intensity but the lowest performance.
Equation 40 below is used to calculate memory bandwidth (BW), where no. of BRAMs $=$ 340, BRAM’s width $=$ 36 KB, no. of LUTRAMs $=$ 129101, LUTRAM’s Width $=$ 32 KB. Our previous work (protea, ), mentioned in the cover letter, provides latency and throughput data (53 GOP/s) in Table 1.
(40) $\begin{split}Memory\ Bandwidth=&(No.\ of\ BRAMs\times BRAM^{\prime}s\ Width+No.\ of\ LUTRAMs\times LUTRAM^{\prime}s\ Width)\times Frequency\end{split}$
@@
Table 1 compares the performance of our accelerator, ADAPTOR, with other FPGA-based accelerators. Each of these accelerators is optimized for specific TNN models, with some designed specifically for sparse computations. TRAC (plagwitz_trac_2022, ) is the only one that automatically generates accelerator code based on the target FPGA and TNN architecture. Since ADAPTOR was synthesized once with fixed hardware resources and bit width, and implemented on a different platform without sparsity, we evaluated metrics such as throughput (GOPS), power consumption, normalized throughput (GOPS per DSP or GOPS per LUT), and power efficiency (GOPS per watt) for a fair comparison. ADAPTOR achieved 1.9$\times$ and 2.25$\times$ higher GOPS compared to the accelerators by Qi et al. in (qi_accommodating_2021, ) and (qi_accelerating_2021, ), respectively, for a shallow transformer. Its normalized throughput was also higher, indicating more efficient DSP and LUT usage without relying on sparsity, whereas Qi et al. applied block balanced pruning and complex block row storage format. Qi et al.’s four-layer transformer encoder (qi_accelerating_2021, ) was 1.7$\times$ slower and 2$\times$ less resource-efficient than ADAPTOR although they applied hierarchical pruning. TRAC (plagwitz_trac_2022, ) used fewer DSPs and LUTs but reported 3.2$\times$ higher GOPS and 8.4$\times$ higher GOPS/DSP. None of these accelerators applied tiling or described any partitioning scheme to accelerate large models such as BERT, but our work includes those. Tzanos et al.’s work (tzanos_hardware_2022, ) is more comparable to ADAPTOR, as they applied tiling and utilized more resources to achieve speeds 1.6 times faster with GOPS /DSP similar to ADAPTOR.